OpenStudy (anonymous):

What is 4.12 expressed as the quotient of two integers in simplest form? Ill Medal :)

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: 4.12 = 4+0.12 = 4 + (12/100)

jimthompson5910 (jim_thompson5910):

you can think of 4 as 400/100

OpenStudy (anonymous):

i was gone the day my class tech this.......i have no idea what this is...can you walk me through?

jimthompson5910 (jim_thompson5910):

do you see how I did 4.12 = 4+0.12 = 4 + (12/100) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

so the answer is 103/25?

jimthompson5910 (jim_thompson5910):

4 + (12/100) = 400/100 + 12/100 = ???

OpenStudy (anonymous):

412/100?

jimthompson5910 (jim_thompson5910):

you can now reduce that

jimthompson5910 (jim_thompson5910):

divide each part by the GCF 4

OpenStudy (anonymous):

so its 100/25

OpenStudy (anonymous):

103 not 100

jimthompson5910 (jim_thompson5910):

103/25, yep

OpenStudy (anonymous):

ok thanks man

OpenStudy (anonymous):

or gal

OpenStudy (anonymous):

im ur fan now :)